OpenStudy (anonymous):

determine whether the ordered pair is a solution of the given equation? (-5,4) x-y=1

myininaya (myininaya):

x=-5 y=4 so replace in the equation x-y=1 -5-4 but this doesn't equal 1 so this is not a ordered pair of the line x-y=1

OpenStudy (anonymous):

I thought so, but your answer help me see that thanks!

myininaya (myininaya):

np